# Disentangled Interpretable Representation for Efficient Long-term Time Series Forecasting

### v1

@@ IV-C1 Basic Components
IV-C2 Low-rank weight sharing
In multivariate time series, our model must effectively capture the relationships between variables. We adopt a low-rank weight sharing approach, specifying the learning of sets of weights. Figure 7 presents the results of testing different values of on the Weather dataset with a prediction length of . When , our model is equivalent to weight sharing, meaning all variables use the same set of weights. When , our model learns a separate set of weights for each variable, treating the variables as independent. The experiments indicate that selecting an appropriate value for enables the model to achieve optimal performance.
We conducted an in-depth analysis to evaluate the necessity of low-rank weight sharing for each component of the model. Extensive experiments were performed on the ETTh2 and Weather datasets, covering multiple prediction horizons. For each component, we compared three configurations: shared weights, low-rank shared weights, and full-rank independent weights. As presented in Table VI, the low-rank weight sharing consistently delivered superior performance across all settings, highlighting its critical role in optimizing model efficacy.
| Components | ETTh2 | Weather | ||||||||
| SFA | STA | IFM | 96 | 192 | 336 | 720 | 96 | 192 | 336 | 720 |
| ✗ | ✓ | ✓ | 0.280 | 0.329 | 0.353 | 0.375 | 0.143 | 0.187 | 0.238 | 0.310 |
| ✓ | ✗ | ✓ | 0.279 | 0.326 | 0.354 | 0.374 | 0.142 | 0.187 | 0.238 | 0.309 |
| ✓ | ✓ | ✗ | 0.283 | 0.329 | 0.353 | 0.378 | 0.142 | 0.188 | 0.238 | 0.310 |
| F | F | F | 0.270 | 0.333 | 0.356 | 0.380 | 0.169 | 0.214 | 0.276 | 0.340 |
| ✓ | ✓ | ✓ | 0.275 | 0.325 | 0.350 | 0.375 | 0.142 | 0.187 | 0.234 | 0.306 |
IV-C3 SFALoss
The parameter in SFALoss governs the trade-off between the frequency domain loss and the time domain loss. To assess the influence of on model performance, we conducted experiments on the ETTh1 and ETTh2 datasets with a prediction length of . The results, illustrated in Figure 8, demonstrate the significant impact of on optimizing the model’s predictive accuracy across both datasets.
IV-D Interpretability: Case Studies
Benefiting from the disentangled representation of our model in both the frequency and time domains, our approach demonstrates superior interpretability. This interpretability facilitates user understanding of the model’s behavior, thereby enhancing trust in its predictive outcomes.
IV-D1 ETTh1
We select the ETTh1 dataset as a simple example. We choose an input length and a prediction horizon , ensuring the model has the minimum number of parameters. Additionally, within the ETTh1 dataset, we set the hyperparameter , which eliminates the router . Thus for each module, we learn only a single set of weights. The weight visualization is shown in Fig. 9. The SFA module explicitly preserves low-frequency information while emphasizing harmonic amplitudes in the high-frequency region. The STA module identifies the importance of each time point in the input sequence for future predictions. The IFM module learns distinct prediction patterns for harmonic frequencies (shown in red) and non-harmonic frequencies (shown in blue).
IV-D2 Weather
In the Weather dataset, different variables represent various meteorological data, such as temperature, pressure, humidity, wind direction, precipitation, etc.. We set as the hyperparameter, allowing the model to learn that the router can represent the correlations among different meteorological data variation patterns. The visualization is shown in Fig. 10.
We can further process the weights learned by the model to understand the knowledge it has captured. Taking the Router as an example, for each variable, the router weight can be viewed as the distribution of its predictive patterns across different weight groups. Consequently, we can compute the Jensen-Shannon Distance (JSD) between variables:
where are two variables, is a mixture distribution, and is the Kullback–Leibler (KL) divergence defined by 17:
Thus, we obtained the JSD Matrix based on the learnt router, as shown in Fig. 11.
IV-D3 Electricity
We further selected the Electricity dataset to comprehensively demonstrate the representations learned by DiPE-Linear. We chose an input length and a prediction horizon in the LSTF setting. In the Electricity dataset, the hyperparameter , meaning that four sets of weights are learned for each module. The parameter visualization is shown in Figure 10. Similar to the ETTh1 dataset, the SFA module learns to retain low-frequency information and high-frequency harmonics. The STA module clearly distinguishes the periodic variations between weekdays and weekends within a month. In the IFM module, four different frequency transformation patterns are also learned. Finally, the weight visualization of the Router illustrates the differences and similarities in predictive patterns across different variables in the dataset.
IV-E Complexity analysis
Our model retains the linear characteristics of FCs while achieving significant reductions in both parameter and computational complexity. In this section, we analyze the complexity of each module in our model in detail. For simplicity, we denote ambiguously both the input length and forecast horizon as .
IV-E1 Parameter Complexity
IV-E2 Computational Complexity
The computational cost of our model primarily arises from two components. First, the model’s parameter computations involve only element-wise multiplications, yielding a computational complexity of . Second, the rFFT and irFFT operations each have a computational complexity of . Therefore, the overall computational complexity of the model is .
A comparison of the parameter and computational complexities between our model and FCs (both with and without weight sharing) is provided in Table VII.
@@ IV-E2 Computational Complexity
| Param. | |||
| Compu. |
V Conclusion
This paper presents DiPE-Linear, a novel approach designed for LTSF in industrial applications. DiPE-Linear effectively addresses the challenges of Industry 5.0, including high-dimensional and high-resolution data as well as high-stakes scenarios, by leveraging disentangled, linear-complexity interpretable modules in both the time and frequency domains. The sophisticated design reduces parameter complexity from quadratic to linear and computational complexity from quadratic to log-linear. Extensive experimental results demonstrate that DiPE-Linear consistently outperforms existing models, significantly enhancing scalability, computational efficiency, and predictive accuracy, particularly in complex industrial environments. In future work, we will enhance DiPE-Linear by incorporating interpretable nonlinear modules, while further exploring lightweight and efficient practical LSTF models.

### v2

@@ 4.3 Ablation Studies
4.3.2 Hyperparameter Analysis and Guidelines.
We analyze the sensitivity of the two key hyperparameters, and , to provide practical tuning guidelines and address concerns about model complexity.
For the Low-rank weight sharing factor , this hyperparameter can be intuitively understood as the number of distinct patterns among the variables in the multivariate time series. Fig. 6a presents the results of varying on the Weather dataset (). The key observation is that model performance remains stable and near-optimal across a wide range of (from 2 to 16). Performance only degrades at the extremes: (full weight sharing) and (full channel independence). This demonstrates that is not a sensitive hyperparameter within a reasonable range, significantly easing the tuning process. In practice, we suggest starting and increasing it exponentially(as ) until validation performance saturates.
For the SFALoss balance parameter , which governs the trade-off between the time-domain loss () and the frequency-domain loss (), we performed a grid search on the ETTh datasets (results in Fig. 6). Our experiments show that the model’s accuracy often has a monotonic or U-shaped relationship with . Notably, the optimal performance is typically found when is in the ranges of [0, 0.3] or [0.7, 1.0]. This suggests the model prefers to optimize primarily for one domain (either time or frequency) rather than an equal mixture. This finding simplifies tuning, as the search space can be effectively concentrated on values near 0 and 1, rather than requiring a fine-grained search across the entire [0, 1] interval.
4.4 Complexity Analysis
Our model retains the linear parameterization of FC models while significantly reducing both parameter and computational complexity. The parameter complexity in the SFA, STA, and IFM modules is . Low-rank weight sharing introduces additional routing parameters with a complexity of , where is the number of channels; this is a higher-order infinitesimal relative to . Thus, the overall parameter complexity remains . For computational complexity, the model’s main costs arise from element-wise multiplications () and the rFFT/irFFT operations (), resulting in a total computational complexity of . Table 4 compares these complexities with those of standard FC models, demonstrating the efficiency of our approach.
We note that while our theoretical computational complexity is superior, practical latency depends on hardware optimizations; FC models benefit from highly optimized matrix multiplication (e.g. cuBLAS), which may not be true for FFT operations. Therefore, the primary advantages of DiPE-Linear are its linear parameter complexity () and drastically reduced memory footprint. This makes DiPE-Linear ideal for resource-constrained environments and for mitigating overfitting on long sequences.
@@ 4.4 Complexity Analysis
| Parameter | |||
| Time |
5 Conclusion
This paper introduces DiPE-Linear, a novel model specifically tailored for LTSF. DiPE-Linear addresses the intrinsic quadratic parameter redundancy and entanglement of existing monolithic FC models by employing disentangled modules—factorizing the modeling task into specialized components for frequential filtering, temporal weighting, and frequential mapping. This innovative architecture, which serves as a strong inductive bias, reduces parameter complexity from quadratic to linear and computational complexity from quadratic to log-linear, significantly improving efficiency and scalability. Comprehensive experimental evaluations demonstrate that DiPE-Linear consistently achieves or surpasses state-of-the-art models, suggesting that for LTSF benchmarks, its well-structured linear design is more effective for generalization than the complex non-linear mapping of larger models. DiPE-Linear thus sets a new baseline for parameter-efficient LTSF models. Future research will explore the integration of nonlinear architectures and further optimization of efficient LTSF frameworks.